# RooflineBench: A Benchmarking Framework for On-Device LLMs via Roofline Analysis

### v1

@@ Abstract
The transition toward localized intelligence through Small Language Models (SLMs) has intensified the need for rigorous performance characterization on resource-constrained edge hardware. However, objectively measuring the theoretical performance ceilings of diverse architectures across heterogeneous platforms remains a formidable challenge. In this work, we propose a systematic framework based on the Roofline model that unifies architectural primitives and hardware constraints through the lens of operational intensity ($OI$). By defining an inference-potential region, we introduce the Relative Inference Potential as a novel metric to compare efficiency differences between Large Language Models (LLMs) on the same hardware substrate. Extensive empirical analysis across diverse compute tiers reveals that variations in performance and $OI$ are significantly influenced by sequence length. We further identify a critical regression in $OI$ as model depth increases. Additionally, our findings highlight an efficiency trap induced by hardware heterogeneity and demonstrate how structural refinements, such as Multi-head Latent Attention ($MLA$), can effectively unlock latent inference potential across various hardware substrates. These insights provide actionable directions for hardware-software co-design to align neural structures with physical constraints in on-device intelligence. The released code is available in the Appendix C.
@@ 1 Introduction
Recent literature has increasingly prioritized the quantitative assessment of inference efficiency through specialized metrics such as Model Bandwidth Utilization ($MBU$) (Agarwal et al., 2023), Model FLOPs Utilization ($MFU$) (Chowdhery et al., 2023), and evaluation frameworks like MoeCap (Jiang et al., 2025). Despite these advancements, establishing a comprehensive understanding of on-device intelligence remains a formidable task. Primarily, objectively characterizing the theoretical performance upper bound for specific large model architectures when deployed on heterogeneous hardware platforms is challenging because of the complex interplay between software kernels and hardware substrates (Dao et al., 2022). Furthermore, conventional evaluation methods often lack the analytical depth required to pinpoint the fundamental physical constraints limiting inference efficacy in resource-constrained environments. These limitations necessitate a more robust analytical framework that can decompose hardware-software interactions to identify precise bottlenecks across diverse compute tiers.
@@
To address these challenges, we leverage the Roofline model as a rigorous analytical bridge to characterize the interplay between large model inference and underlying hardware capabilities (Williams et al., 2009). Our approach entails the empirical measurement of peak $FLOPS$ and memory bandwidth to establish a realistic performance envelope across diverse platforms. We also argue that LLMs exhibit an inference potential region on edge hardware devices (the blue region in in Figure 1), and we propose a novel metric (Relative Inference Potential) to compare the differences in inference potential between two LLMs on the same hardware device. The primary contributions of this work are three fold:
  • Integrated Benchmarking Framework: We propose a systematic Roofline based framework that unifies architectural primitives and hardware constraints via operational intensity ($OI$), defining an inference potential region to introduce the Relative Inference Potential ($\Phi$) for comparative efficiency analysis.
  • Comprehensive Empirical Analysis: Through extensive experiments across heterogeneous compute tiers, we reveal that inference efficiency is primarily governed by context length and attention architectures, while identifying a critical $OI$ regression at increased model depths.
@@ 2 Background
For a decoder-only model with $n_{params}$ parameters, it has been shown that the forward pass requires approximately $2\cdot n_{params}$ floating-point operations (FLOPs) per token . Let $P_{peak}$ be the peak performance of the hardware and $BW$ be the memory bandwidth. During decoding, the model weights and the attention key-value (KV) cache must be transferred from memory to compute cores for every single token. We define the compute time ($T_{comp}$) and memory time ($T_{mem}$) as:
@@
On resource-constrained edge devices, the additional memory traffic from the KV cache further reduces the system’s Operational Intensity (OI), shifting the execution point to the left on the Roofline model (Williams et al., 2009). Simultaneously, the limited memory bandwidth ($BW$) of mobile SoCs causes the hardware’s “knee point” ($P_{peak}/BW$) to shift to the right. This dual effect forces the decoding phase deep into the severe memory-bandwidth bound region, where the computational cores remain underutilized while waiting for tensor loading. Consequently, architectural optimizations such as multiquery attention (MQA) are essential to reduce the KV cache footprint and alleviate this bandwidth pressure.
@@ 3.1 Standard Roofline Formulation
Where $P_{peak}$: Theoretical peak compute performance (GFLOPS). $BW_{peak}$: Achievable peak memory bandwidth (GB/s). $OI$ (Operational Intensity): The ratio of floating-point operations to bytes of memory traffic (FLOPs/Byte). Given the autoregressive nature of LLM decoding (token-by-token generation), the process is inherently memory-bound. Consequently, our analysis focuses primarily on the sloped region of the Roofline graph, where performance is strictly limited by $OI\times BW_{peak}$.
@@ 3.2 Applying the Standard Roofline Model to LLMs
  • FLOPs ($W$): Analytical Estimation. Directly measuring FLOPs via hardware counters can be imprecise due to instrumentation overhead. Instead, we adopt the analytical formulation from Sun et al. (2025) to estimate the theoretical FLOPs for heterogeneous architectures (e.g., MHA, GQA). Given a model with hidden dimension $H$, sequence length $N$, and head configurations $\{n_{q},n_{k},n_{v}\}$, we precisely calculate the computational load for the Linear, Attention, and FFN layers per decoding step. This ensures a uniform standard for calculating Performance ($P$) and Operational Intensity ($OI$) across different backends.
@@ 3.3 Characterizing Relative Inference Potential across Performance Regimes
Compute Bound Regime ($OI\geq OI_{r}$): Since horizontal $OI$ gains yield negligible ceiling increases, $\Phi$ is defined as the vertical distance to the peak compute limit $\pi$:
@@ 4 Comprehensive Characterization Study
This section provides a systematic characterization of LLM inference using the proposed Relative Inference Potential ($\Phi$) within a Roofline analytical framework. Unlike traditional macro metrics such as throughput (TPS), $\Phi$ facilitates a regime aware evaluation of execution efficiency by quantifying the spatial discrepancy between realized performance and theoretical hardware limits. We evaluate task level sensitivity across four sequence patterns (Sec. 4.1), investigate parameter level scaling and bottleneck shifts through model depth profiling (Sec. 4.2), and assess algorithmic impacts from precision and architectural refinements such as $MLA$ (Sec. 4.3). By identifying critical transitions between memory bound and compute bound regimes, this multidimensional analysis leverages $\Phi$ to reveal the internal complexity and optimization headroom of localized intelligence. Experimental details are in Appendix C.
@@ 4.1 Sensitivity to Input-Output Sequence Lengths
In Figure 2, decoding performance exhibits significant sensitivity to sequence lengths across all tested models. Notably, the LISO scenario (yellow dots) consistently achieves the highest execution efficiency, positioning itself as the closest observed case to the Roofline ridge point. This proximity stems from the larger input context, which increases the relative computational demand of the attention mechanism within each decoding step. By effectively amortizing the fixed memory overhead of loading model weights, LISO exhibits a high computational proportion, thereby elevating both the operational intensity ($OI$) and attainable performance ($GFLOPS$) toward the hardware’s compute-bound limit. Conversely, the SILO scenario (red dots) resides deep within the memory-bound regime. In this scenario, the negligible computational requirement of processing a minimal context cannot offset the massive data movement of weights, leading to severe hardware underutilization.
@@ 4.2 Bottleneck Evolution under Parameter Expansion
As illustrated in Figure 3, scaling model depth reveals a distinct, non-linear evolution of hardware utilization on edge platforms. Initially, the attainable performance ($GFLOPS$) rises consistently and becomes increasingly dense as layers increase (transitioning from light yellow to dark purple dots). This behavior suggests that deeper models more effectively amortize fixed system-level overheads—such as kernel launch latencies, memory synchronization, and command orchestration—thereby enabling the hardware to operate closer to its steady-state compute limits.
More significantly, the operational intensity ($OI$) exhibits a non-monotonic ”arch” trajectory with a clear inflection point at remarkably shallow depths. As depth increases from 2 to approximately 3–5 layers, the data points initially shift to the right, indicating an initial gain in arithmetic reuse as the system moves away from the overhead-dominated regime. However, beyond this threshold, the $OI$ begins to retreat to the left. In resource-constrained on-device environments, this regression arises because the cumulative memory bandwidth pressure of streaming weights for additional layers begins to outpace the marginal gains in computational reuse during the decoding phase. Consequently, the hardware ”memory-wall” is encountered significantly earlier than theoretical predictions suggest, forcing the decoding process into an increasingly memory-bound regime as the model scales, which further constrains the release of relative inference-potential.
@@ 4.3.1 Numerical Precision: From FP16 to 4-bit Quantization
In Figure 4, lowering numerical precision triggers a diagonal shift of the data points toward the upper-right region of the Roofline plane. However, the magnitude of this improvement is highly sensitive to the task-level bottleneck. In memory-bound scenarios such as SILO (bottom-left), transitioning from FP16 to Q4_K_M yields dramatic gains in both operational intensity (OI) and attainable performance ($GFLOPS$), as the reduced bandwidth pressure directly accelerates the execution. Conversely, in the LISO scenario (top-right), the clusters for the three precisions appear significantly tighter. Since LISO is already positioned near the Roofline ridge with high arithmetic intensity, the benefits of reduced data movement are less pronounced, indicating that the execution has largely transitioned from being memory-constrained to being limited by the hardware’s peak compute capability.
@@
In Figure 6, the choice of attention mechanism significantly shifts the inference profile on the Roofline plane. MLA (blue dots) consistently achieves the highest operational intensity ($OI$) and attainable performance ($GFLOPS$) across all four sequence scenarios. By utilizing latent compression for the KV cache, MLA substantially reduces the volume of data movement required for each decoding step compared to the standard MHA, thereby shifting the execution significantly closer to the Roofline ridge point. Interestingly, at this specific 1.5B scale on edge hardware, GQA (red dots) exhibits the lowest efficiency among the three. It indicates that while GQA reduces the number of Key/Value heads, the latent-based compression in MLA provides a more effective balance between memory traffic reduction and computational throughput for on-device decoding.
@@ 5.1 Disparity across Heterogeneous Hardware Platforms
The critical indicator of hardware unfairness is the theoretical ridge point, which dictates the minimum $OI$ required to saturate computational resources. High-performance GPUs such as the RTX 3090 (Ridge: 38.00) and RTX 3070Ti Laptop (Ridge: 37.05) possess significantly higher ridge points, meaning they require a much higher $OI$ to transition from memory-bound to compute-bound regimes. In contrast, edge-side chips like the Apple M1 Pro (Ridge: 25.39), Jetson Orin Nano (Ridge: 20.48) and Raspberry Pi 5 (Ridge: 8.98) have much lower ridge points. While these devices are easier to saturate, their absolute performance ceilings are considerably lower. This disparity implies that a specific model architecture, such as a 1.5B model in the LISO scenario, may achieve near-optimal saturation on a Jetson Orin or Raspberry Pi but remain severely underutilized on an RTX 3090 because of the structural differences in theoretical memory-to-compute ratios.
@@ 5.2 Architectural Sensitivity of Utilization Equity
Despite the collective performance gains, the relative hierarchy of utilization equity remains sensitive to the specific architectural design. MLA (blue markers) consistently maintains a superior position in both operational intensity ($OI$) and attainable performance ($GFLOPS$) on both platforms. Rather than being a platform-specific enhancement, the latent-based compression in MLA functions as a robust efficiency baseline. This stability proves that optimized architectural structures can reliably capture available hardware resources across the device spectrum, ensuring that the performance advantages of high-end accelerators are not overshadowed by the structural inefficiencies of traditional attention mechanisms.
@@ 6 Discussion
The architectural evolution toward Multi head Latent Attention (MLA), as seen in DeepSeek V2(DeepSeek-AI, 2024) and PLM(Deng et al., 2025), along with the trainable sparse attention mechanism in MiniCPM(Hu et al., 2024), exemplifies the necessity of hardware aware innovation to transcend the memory wall in on device environments. MLA utilizes latent compression for the KV cache to reduce the volume of data movement required for each decoding step. Simultaneously, the sparse attention in MiniCPM reduces computational overhead by processing less than $5\%$ of tokens in long context scenarios. Both structural choices elevate the operational intensity ($OI$) and shift the inference profile toward the compute bound regime. As shown in the cross platform comparison, MLA based configurations achieve a higher proficiency in capturing available computational peaks on high performance hardware compared to traditional mechanisms. Such synergy suggests that future on device architectures should prioritize optimizations such as latent compression or trainable sparsity to better align with the divergent theoretical ridge points of heterogeneous hardware.
@@
Capacity density, defined as the ratio of effective to actual parameter size, provides a unified framework for evaluating model quality according to the Densing Law(Xiao et al., 2024). This metric is crucial for resource-constrained environments where traditional scaling through model size is increasingly unsustainable. Layer-wise analysis of models such as Qwen2.5(Yang et al., 2024) reveals that operational intensity ($OI$) does not scale linearly with depth but reaches a peak at low layer counts. In the LISO scenario, increasing depth beyond this optimal threshold leads to an $OI$ fallback. Because edge platforms such as the Raspberry Pi 5 or Jetson Orin Nano have low theoretical ridge points, they are highly sensitive to $OI$ fluctuations. Consequently, improving the capacity density of shallow architectures is a more sustainable strategy for on-device intelligence than simply stacking layers, as it maximizes effective parameters within a restricted memory footprint.
@@
The gap between theoretical peaks and realized throughput suggests that framework design must prioritize data movement over raw compute capacity. Although $TFLOPS$ have grown rapidly, inference remains fundamentally limited by memory access patterns across diverse devices. Refining the interaction between hierarchical memory structures and execution kernels is essential to mitigate the bandwidth constraints that govern large scale model deployment. Additionally, hardware heterogeneity necessitates specializing AI compute units for specific high load primitives. General purpose logic frequently fails to exploit the structural advantages of modern architectures, leading to efficiency losses in on device scenarios. Hardware level optimization for critical primitives such as MLA and sparse attention is therefore crucial for maximizing throughput. Dedicated silicon support ensures that the theoretical operational intensity ($OI$) of optimized models can reach its full potential in real world execution.

### v2

# arXiv:2602.11506v2[cs.LG] 27 Feb 2026

@@ Abstract
The transition toward localized intelligence through Small Language Models (SLMs) has intensified the need for rigorous performance characterization on resource-constrained edge hardware. However, objectively measuring the theoretical performance ceilings of diverse architectures across heterogeneous platforms remains a formidable challenge. In this work, we propose a systematic framework based on the Roofline model that unifies architectural primitives and hardware constraints through the lens of operational intensity (OIOI). By defining an inference-potential region, we introduce the Relative Inference Potential as a novel metric to compare efficiency differences between Large Language Models (LLMs) on the same hardware substrate. Extensive empirical analysis across diverse compute tiers reveals that variations in performance and OIOI are significantly influenced by sequence length. We further identify a critical regression in OIOI as model depth increases. Additionally, our findings highlight an efficiency trap induced by hardware heterogeneity and demonstrate how structural refinements, such as Multi-head Latent Attention (MLAMLA), can effectively unlock latent inference potential across various hardware substrates. These insights provide actionable directions for hardware-software co-design to align neural structures with physical constraints in on-device intelligence. The released code is available in the Appendix C.
@@ 1 Introduction
Recent literature has increasingly prioritized the quantitative assessment of inference efficiency through specialized metrics such as Model Bandwidth Utilization (MBUMBU) (Agarwal et al., 2023), Model FLOPs Utilization (MFUMFU) (Chowdhery et al., 2023), and evaluation frameworks like MoeCap (Jiang et al., 2025). Despite these advancements, establishing a comprehensive understanding of on-device intelligence remains a formidable task. Primarily, objectively characterizing the theoretical performance upper bound for specific large model architectures when deployed on heterogeneous hardware platforms is challenging because of the complex interplay between software kernels and hardware substrates (Dao et al., 2022). Furthermore, conventional evaluation methods often lack the analytical depth required to pinpoint the fundamental physical constraints limiting inference efficacy in resource-constrained environments. These limitations necessitate a more robust analytical framework that can decompose hardware-software interactions to identify precise bottlenecks across diverse compute tiers.
@@
To address these challenges, we leverage the Roofline model as a rigorous analytical bridge to characterize the interplay between large model inference and underlying hardware capabilities (Williams et al., 2009). Our approach entails the empirical measurement of peak FLOPSFLOPS and memory bandwidth to establish a realistic performance envelope across diverse platforms. We also argue that LLMs exhibit an inference potential region on edge hardware devices (the blue region in in Figure 1), and we propose a novel metric (Relative Inference Potential) to compare the differences in inference potential between two LLMs on the same hardware device. The primary contributions of this work are three fold:
  • Integrated Benchmarking Framework: We propose a systematic Roofline based framework that unifies architectural primitives and hardware constraints via operational intensity (OIOI), defining an inference potential region to introduce the Relative Inference Potential ($\Phi$) for comparative efficiency analysis.
  • Comprehensive Empirical Analysis: Through extensive experiments across heterogeneous compute tiers, we reveal that inference efficiency is primarily governed by context length and attention architectures, while identifying a critical OIOI regression at increased model depths.
@@ 2 Background
For a decoder-only model with $n_{params}$ parameters, it has been shown that the forward pass requires approximately $2\cdot n_{params}$ floating-point operations (FLOPs) per token . Let $P_{peak}$ be the peak performance of the hardware and BWBW be the memory bandwidth. During decoding, the model weights and the attention key-value (KV) cache must be transferred from memory to compute cores for every single token. We define the compute time ($T_{comp}$) and memory time ($T_{mem}$) as:
@@
On resource-constrained edge devices, the additional memory traffic from the KV cache further reduces the system’s Operational Intensity (OI), shifting the execution point to the left on the Roofline model (Williams et al., 2009). Simultaneously, the limited memory bandwidth (BWBW) of mobile SoCs causes the hardware’s “knee point” ($P_{peak}/BW$) to shift to the right. This dual effect forces the decoding phase deep into the severe memory-bandwidth bound region, where the computational cores remain underutilized while waiting for tensor loading. Consequently, architectural optimizations such as multiquery attention (MQA) are essential to reduce the KV cache footprint and alleviate this bandwidth pressure.
@@ 3.1 Standard Roofline Formulation
Where $P_{peak}$: Theoretical peak compute performance (GFLOPS). $BW_{peak}$: Achievable peak memory bandwidth (GB/s). OIOI (Operational Intensity): The ratio of floating-point operations to bytes of memory traffic (FLOPs/Byte). Given the autoregressive nature of LLM decoding (token-by-token generation), the process is inherently memory-bound. Consequently, our analysis focuses primarily on the sloped region of the Roofline graph, where performance is strictly limited by $OI\times BW_{peak}$.
@@ 3.2 Applying the Standard Roofline Model to LLMs
  • FLOPs ($W$): Analytical Estimation. Directly measuring FLOPs via hardware counters can be imprecise due to instrumentation overhead. Instead, we adopt the analytical formulation from Sun et al. (2025) to estimate the theoretical FLOPs for heterogeneous architectures (e.g., MHA, GQA). Given a model with hidden dimension $H$, sequence length NN, and head configurations $\{n_{q},n_{k},n_{v}\}$, we precisely calculate the computational load for the Linear, Attention, and FFN layers per decoding step. This ensures a uniform standard for calculating Performance ($P$) and Operational Intensity (OIOI) across different backends.
@@ 3.3 Characterizing Relative Inference Potential across Performance Regimes
Compute Bound Regime ($OI\geq OI_{r}$): Since horizontal OIOI gains yield negligible ceiling increases, $\Phi$ is defined as the vertical distance to the peak compute limit $\pi$:
@@ 4 Comprehensive Characterization Study
This section provides a systematic characterization of LLM inference using the proposed Relative Inference Potential ($\Phi$) within a Roofline analytical framework. Unlike traditional macro metrics such as throughput (TPS), $\Phi$ facilitates a regime aware evaluation of execution efficiency by quantifying the spatial discrepancy between realized performance and theoretical hardware limits. We evaluate task level sensitivity across four sequence patterns (Sec. 4.1), investigate parameter level scaling and bottleneck shifts through model depth profiling (Sec. 4.2), and assess algorithmic impacts from precision and architectural refinements such as MLAMLA (Sec. 4.3). By identifying critical transitions between memory bound and compute bound regimes, this multidimensional analysis leverages $\Phi$ to reveal the internal complexity and optimization headroom of localized intelligence. Experimental details are in Appendix C.
@@ 4.1 Sensitivity to Input-Output Sequence Lengths
In Figure 2, decoding performance exhibits significant sensitivity to sequence lengths across all tested models. Notably, the LISO scenario (yellow dots) consistently achieves the highest execution efficiency, positioning itself as the closest observed case to the Roofline ridge point. This proximity stems from the larger input context, which increases the relative computational demand of the attention mechanism within each decoding step. By effectively amortizing the fixed memory overhead of loading model weights, LISO exhibits a high computational proportion, thereby elevating both the operational intensity (OIOI) and attainable performance (GFLOPSGFLOPS) toward the hardware’s compute-bound limit. Conversely, the SILO scenario (red dots) resides deep within the memory-bound regime. In this scenario, the negligible computational requirement of processing a minimal context cannot offset the massive data movement of weights, leading to severe hardware underutilization.
@@ 4.2 Bottleneck Evolution under Parameter Expansion
As illustrated in Figure 3, scaling model depth reveals a distinct, non-linear evolution of hardware utilization on edge platforms. Initially, the attainable performance (GFLOPSGFLOPS) rises consistently and becomes increasingly dense as layers increase (transitioning from light yellow to dark purple dots). This behavior suggests that deeper models more effectively amortize fixed system-level overheads—such as kernel launch latencies, memory synchronization, and command orchestration—thereby enabling the hardware to operate closer to its steady-state compute limits.
More significantly, the operational intensity (OIOI) exhibits a non-monotonic ”arch” trajectory with a clear inflection point at remarkably shallow depths. As depth increases from 2 to approximately 3–5 layers, the data points initially shift to the right, indicating an initial gain in arithmetic reuse as the system moves away from the overhead-dominated regime. However, beyond this threshold, the OIOI begins to retreat to the left. In resource-constrained on-device environments, this regression arises because the cumulative memory bandwidth pressure of streaming weights for additional layers begins to outpace the marginal gains in computational reuse during the decoding phase. Consequently, the hardware ”memory-wall” is encountered significantly earlier than theoretical predictions suggest, forcing the decoding process into an increasingly memory-bound regime as the model scales, which further constrains the release of relative inference-potential.
@@ 4.3.1 Numerical Precision: From FP16 to 4-bit Quantization
In Figure 4, lowering numerical precision triggers a diagonal shift of the data points toward the upper-right region of the Roofline plane. However, the magnitude of this improvement is highly sensitive to the task-level bottleneck. In memory-bound scenarios such as SILO (bottom-left), transitioning from FP16 to Q4_K_M yields dramatic gains in both operational intensity (OI) and attainable performance (GFLOPSGFLOPS), as the reduced bandwidth pressure directly accelerates the execution. Conversely, in the LISO scenario (top-right), the clusters for the three precisions appear significantly tighter. Since LISO is already positioned near the Roofline ridge with high arithmetic intensity, the benefits of reduced data movement are less pronounced, indicating that the execution has largely transitioned from being memory-constrained to being limited by the hardware’s peak compute capability.
@@
In Figure 6, the choice of attention mechanism significantly shifts the inference profile on the Roofline plane. MLA (blue dots) consistently achieves the highest operational intensity (OIOI) and attainable performance (GFLOPSGFLOPS) across all four sequence scenarios. By utilizing latent compression for the KV cache, MLA substantially reduces the volume of data movement required for each decoding step compared to the standard MHA, thereby shifting the execution significantly closer to the Roofline ridge point. Interestingly, at this specific 1.5B scale on edge hardware, GQA (red dots) exhibits the lowest efficiency among the three. It indicates that while GQA reduces the number of Key/Value heads, the latent-based compression in MLA provides a more effective balance between memory traffic reduction and computational throughput for on-device decoding.
@@ 5.1 Disparity across Heterogeneous Hardware Platforms
The critical indicator of hardware unfairness is the theoretical ridge point, which dictates the minimum OIOI required to saturate computational resources. High-performance GPUs such as the RTX 3090 (Ridge: 38.00) and RTX 3070Ti Laptop (Ridge: 37.05) possess significantly higher ridge points, meaning they require a much higher OIOI to transition from memory-bound to compute-bound regimes. In contrast, edge-side chips like the Apple M1 Pro (Ridge: 25.39), Jetson Orin Nano (Ridge: 20.48) and Raspberry Pi 5 (Ridge: 8.98) have much lower ridge points. While these devices are easier to saturate, their absolute performance ceilings are considerably lower. This disparity implies that a specific model architecture, such as a 1.5B model in the LISO scenario, may achieve near-optimal saturation on a Jetson Orin or Raspberry Pi but remain severely underutilized on an RTX 3090 because of the structural differences in theoretical memory-to-compute ratios.
@@ 5.2 Architectural Sensitivity of Utilization Equity
Despite the collective performance gains, the relative hierarchy of utilization equity remains sensitive to the specific architectural design. MLA (blue markers) consistently maintains a superior position in both operational intensity (OIOI) and attainable performance (GFLOPSGFLOPS) on both platforms. Rather than being a platform-specific enhancement, the latent-based compression in MLA functions as a robust efficiency baseline. This stability proves that optimized architectural structures can reliably capture available hardware resources across the device spectrum, ensuring that the performance advantages of high-end accelerators are not overshadowed by the structural inefficiencies of traditional attention mechanisms.
@@ 6 Discussion
The architectural evolution toward Multi head Latent Attention (MLA), as seen in DeepSeek V2(DeepSeek-AI, 2024) and PLM(Deng et al., 2025), along with the trainable sparse attention mechanism in MiniCPM(Hu et al., 2024), exemplifies the necessity of hardware aware innovation to transcend the memory wall in on device environments. MLA utilizes latent compression for the KV cache to reduce the volume of data movement required for each decoding step. Simultaneously, the sparse attention in MiniCPM reduces computational overhead by processing less than $5\%$ of tokens in long context scenarios. Both structural choices elevate the operational intensity (OIOI) and shift the inference profile toward the compute bound regime. As shown in the cross platform comparison, MLA based configurations achieve a higher proficiency in capturing available computational peaks on high performance hardware compared to traditional mechanisms. Such synergy suggests that future on device architectures should prioritize optimizations such as latent compression or trainable sparsity to better align with the divergent theoretical ridge points of heterogeneous hardware.
@@
Capacity density, defined as the ratio of effective to actual parameter size, provides a unified framework for evaluating model quality according to the Densing Law(Xiao et al., 2024). This metric is crucial for resource-constrained environments where traditional scaling through model size is increasingly unsustainable. Layer-wise analysis of models such as Qwen2.5(Yang et al., 2024) reveals that operational intensity (OIOI) does not scale linearly with depth but reaches a peak at low layer counts. In the LISO scenario, increasing depth beyond this optimal threshold leads to an OIOI fallback. Because edge platforms such as the Raspberry Pi 5 or Jetson Orin Nano have low theoretical ridge points, they are highly sensitive to OIOI fluctuations. Consequently, improving the capacity density of shallow architectures is a more sustainable strategy for on-device intelligence than simply stacking layers, as it maximizes effective parameters within a restricted memory footprint.
@@
The gap between theoretical peaks and realized throughput suggests that framework design must prioritize data movement over raw compute capacity. Although TFLOPSTFLOPS have grown rapidly, inference remains fundamentally limited by memory access patterns across diverse devices. Refining the interaction between hierarchical memory structures and execution kernels is essential to mitigate the bandwidth constraints that govern large scale model deployment. Additionally, hardware heterogeneity necessitates specializing AI compute units for specific high load primitives. General purpose logic frequently fails to exploit the structural advantages of modern architectures, leading to efficiency losses in on device scenarios. Hardware level optimization for critical primitives such as MLA and sparse attention is therefore crucial for maximizing throughput. Dedicated silicon support ensures that the theoretical operational intensity (OIOI) of optimized models can reach its full potential in real world execution.